# 3D-GRAND: A Million-Scale Dataset for 3D-LLMs with Better Grounding and Less Hallucination

### v2

@@ Abstract
The integration of language and 3D perception is crucial for developing embodied agents and robots that comprehend and interact with the physical world. While large language models (LLMs) have demonstrated impressive language understanding and generation capabilities, their adaptation to 3D environments (3D-LLMs) remains in its early stages. A primary challenge is the absence of large-scale datasets that provide dense grounding between language and 3D scenes. In this paper, we introduce 3D-GRAND, a pioneering large-scale dataset comprising 40,087 household scenes paired with 6.2 million densely-grounded scene-language instructions. Our results show that instruction tuning with 3D-GRAND significantly enhances grounding capabilities and reduces hallucinations in 3D-LLMs. As part of our contributions, we propose a comprehensive benchmark 3D-POPE to systematically evaluate hallucination in 3D-LLMs, enabling fair comparisons among future models. Our experiments highlight a scaling effect between dataset size and 3D-LLM performance, emphasizing the critical role of large-scale 3D-text datasets in advancing embodied AI research. Notably, our results demonstrate early signals for effective sim-to-real transfer, indicating that models trained on large synthetic data can perform well on real-world 3D scans. Through 3D-GRAND and 3D-POPE, we aim to equip the embodied AI community with essential resources and insights, setting the stage for more reliable and better-grounded 3D-LLMs.
Figure 1: We introduce 3D-GRAND, a large-scale, densely grounded 3D-text dataset, and 3D-POPE, a 3D-LLM hallucination benchmark. Training on 3D-GRAND improves grounding accuracy and reduces hallucinations.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have greatly benefited Embodied Artificial Intelligence (EAI). LLMs demonstrate exceptional capabilities in understanding language instructions [49, 64], perceiving the environment [44, 39, 3, 81, 71], and planning detailed actions [7, 33]. The primary inputs to LLMs, other than pure language, have been the combination of language and 2D images, categorizing these models as 2D-LLMs. The significant advancements in 2D-LLMs can be largely attributed to their training on extensive vision-language datasets. These datasets [58, 82], comprising billions of image and text pairs, have been instrumental in enhancing the models’ understanding of visual content and its contextual relevance to textual information. These large datasets have provided the foundational data necessary for training models that excel at integrating visual perception with language processing. Despite some progress in equipping LLMs to understand 3D scenes (3D-LLMs) [27, 29, 66, 30, 83, 11, 53], these models remain in their early stages due to the scarcity of 3D scene and text pairs. In this work, we introduce 3D-GRAND, a pioneering million-scale dataset designed for densely-grounded 3D Instruction Tuning.
Recently, SceneVerse [35] concurrently introduced a large-scale 3D vision-language dataset. However, a significant limitation of this dataset is the absence of object grounding in language, which is crucial for enhancing model usability in robotics tasks and reducing hallucination. Research on 2D-LLMs indicates that grounding language to 2D contexts notably mitigates hallucination in language models [74, 51, 5, 38, 54, 78], thereby enhancing the reliability and interpretability of generated responses. While 2D grounding has been extensively explored, extending these principles to 3D environments is still underdeveloped. This situation raises two critical questions: (1) Is there any hallucination in 3D-LLMs and if so, how severe is it? (2) Can densely-grounded data mitigate hallucination for 3D-LLMs? These questions underscore a critical need within the research community for the development of an evaluation benchmark specifically designed for 3D-LLMs and the construction of a large-scale, 3D-grounded dataset.
To quantify hallucination in 3D LLMs, this work introduces 3D-POPE (3D Polling-based Object Probing Evaluation). 3D-POPE provides a comprehensive and standardized protocol for evaluating hallucination that enables systematic assessment and facilitates fair comparisons across 3D-LLMs, enhancing our understanding of model capabilities in object hallucination. Specifically, we pose existence questions to 3D-LLMs and evaluate their responses, as illustrated in Fig 1.
To evaluate the role of densely-grounded dataset, we introduce a pioneering million-scale dataset, 3D-GRAND, for densely grounded 3D instruction tuning. 3D-GRAND includes 40,087 household scenes paired with 6.2 million scene-language instructions, featuring dense phrase-to-object grounding. We conduct rigorous human evaluations to ensure the dataset’s quality. Our results trained with 3D-GRAND highlight the dataset’s effectiveness in enhancing grounding and reducing hallucination for 3D-LLMs. We highlight the effectiveness of incorporating 3D-GRAND in Fig 1 and introduce each category of 3D-GRAND and provide examples in Fig 2.
To sum up, our contributions include:
  • 3D-GRAND, the first million-scale, densely-grounded 3D-text dataset for grounded 3D Instruction Tuning. 3D-GRAND includes 40K household scenes paired with 6.2M densely-grounded scene-language instructions.
  • 3D-POPE, a suite of benchmarks and metrics that systematically evaluate hallucination, enabling fair comparisons of future 3D-LLM models in terms of object hallucination.
  • Quantitative research findings regarding hallucination, grounding, and scaling that provide guidance to future research: (1). training 3D-LLMs with 3D-GRAND significantly reduces hallucinations, particularly when the data is densely grounded; (2). densely grounded instruction tuning significantly enhances the grounding capabilities of 3D-LLMs; (3). scaling densely grounded data consistently improves grounding accuracy and reduces hallucination; and (4). models can successfully transfer from sim-to-real, providing an early signal for a low-cost and sustainable future of scaling synthetic 3D data to help on real tasks.
@@
Table 1: Comparison of 3D-GRAND with existing 3D scene datasets with language annotations. 3D-GRAND is the largest language-grounded dataset.
@@ 2 Related Work
Injecting 3D into LLMs. Recent advancements in large language models (LLMs) have inspired research into extending their capabilities to 3D environments, leading to the development of 3D-LLMs [11, 53, 71, 83]. Notable works in this field include 3D-LLM [27], which integrates 3D point clouds and features into LLMs to enable tasks such as captioning, question answering, and navigation. LEO [30] excels as an embodied multi-modal generalist agent in perception, grounding, reasoning, planning, and action in 3D environments, highlighting the potential of 3D-LLMs in understanding and interacting with the physical world. The most relevant work to our model is Chat-3Dv2 [66, 29], which grounds generated scene captions to objects in 3D scenes. However, Chat-3Dv2’s dataset is limited to one type of 3D-text task (scene captioning) and only includes 705 captions from a subset of ScanNet scenes. In 3D-GRAND, we expand this concept by diversifying 3D-text tasks and increasing the dataset size to a million-scale. Our results demonstrate promising data scaling effects and sim-to-real transfer, paving the way for future large-scale training of 3D-LLMs.
@@
Object Hallucination of VLMs. While 2D VLMs have achieved impressive performance, they are prone to hallucinating objects that do not exist in the provided images, a problem known as object hallucination [15, 56]. Several methods have been suggested to mitigate the object hallucination issue, such as integrating an external object detector [77], applying visually grounded visual instruction tuning [74, 78] or reinforcement learning [61, 24], performing iterative refinement [80], and adapting the decoding strategies [31]. To quantify and mitigate this issue, several benchmarks have been proposed. CHAIR (Caption Hallucination Assessment with Image Relevance) [56] measures the frequency of hallucinated objects in image captions by comparing the objects mentioned to the ground truth annotations. POPE (Probing Object Hallucination Evaluation) [41] assesses a VLM’s ability to identify the presence or absence of objects through yes/no probing questions. However, these studies primarily focus on 2D image-text datasets like COCO [43]. In contrast, object hallucination in 3D-LLMs remains largely unexplored. Our work addresses this gap by introducing 3D-POPE, a comprehensive benchmark for evaluating object hallucination in 3D-LLMs. To the best of our knowledge, this is the first object hallucination benchmark for 3D-LLMs.
@@
Grounding datasets. In the 2D domain, large-scale datasets with grounding information have been instrumental in advancing vision-language research. Notable examples include RefCOCO [75], which provides referring expressions for objects in COCO images [43]. Additionally, 2D LLMs [51, 54, 69, 38, 74] have been trained with densely-grounded web-crawled image-text pairs. In the 3D domain, there is a growing interest in creating datasets that pair 3D scenes with textual annotations [76, 1, 32, 12]. ScanRefer [10] pioneered this effort by contributing a dataset of ScanNet [14] scenes with referring expressions. Table 1 introduces the efforts made by the community. However, these datasets have limited grounding annotations and often focus on a single task, such as referring expression comprehension or visual question answering. In contrast, our proposed dataset, 3D-GRAND, stands out by providing 6.2 million densely-grounded scene-language instructions across a diverse set of 3D-text tasks and 40,087 household scenes. This enables a wide range of grounding tasks and facilitates the development of more reliable and better-grounded 3D-LLMs.
@@
3 3D-GRAND: The 3D Ground Anything Dataset
In this section, we introduce 3D-GRAND, a large-scale, densely-grounded 3D-text dataset designed for grounded 3D instruction tuning. We describe the data collection process, dataset statistics, and the unique features that make 3D-GRAND a valuable resource for advancing research in 3D-LLMs.
@@
To this end, we curate a diverse collection of 40,087 high-quality 3D indoor scenes from the 3D-FRONT [23] and Structured3D [79] datasets. These datasets are chosen for their large quantities of synthetic indoor scenes with professionally designed layouts. The collection includes a variety of room types, such as living rooms, bedrooms, kitchens, office spaces, and conference rooms. We further process these 3D scenes to generate per-room 3D point clouds. Details on point cloud rendering and cleaning are provided in the Appendix.
@@
The second, and perhaps more critical, motivation is the scalability of annotation. While we can potentially scale up 3D scenes using synthetic data generation pipelines, annotating these scenes with human effort is both costly and time-consuming, especially for complex tasks like densely grounding annotation. To put the cost of money and time in perspective, for the data we annotated in this paper, we estimate that obtaining the same annotations with human annotator would cost at least \$539,000 and require 5.76 years (no eat, no sleep) worth of work from a professional annotator (earning minimum wage of \$10.67 per hour). In contrast, using LLMs (GPT4 [49]), we achieve the same results for \$3,030 within 2 days, representing a 178x reduction in cost and a 1051x reduction in time. At the time of writing, the cost and time further decreases by 50% to \$1,500 and 1 day, with the introduction of GPT-4o [50].
@@
$\bullet$ 3D Model to 2D Image. In the 3D-Front dataset, each object is sourced from 3D-Future [23], which provides a ground truth 2D image for each object. For the Structured3D dataset, individual images for each object are not available. Therefore, we utilize the set-of-mark prompting technique [72], where each object to be annotated is circled in red in the images.
$\bullet$ 2D Image to Attributes. We use GPT-4V to generate detailed language annotations for each 2D object image, including attributes like name, color, finish, and texture. The naming is now open-vocabulary, contrary to being class-agnostic.
@@
$\bullet$ Scene Graph to Generated Annotations. Based on the given scene graph, we will be able to produce 3D-Grounded Object Reference, 3D-Grounded Scene Description and 3D-Grounded QA using GPT-4 [49] and GPT4-Turbo with various prompts, which we will show in the appendix.
$\bullet$ Generated Annotations to Processed Annotations. After we acquire raw annotations, we will apply hallucination filters and template augmentation for the phrase tag to remove low quality annotations and augment generated annotations.
We want to reinforce the set of grounded language tasks that cover a diverse range of language understanding and generation scenarios in 3D environments which we show in Figure 2:
@@
Dataset highlights. 3D-GRAND possesses several unique features that distinguish it from existing 3D-language datasets: (1). Large-scale: With 40,087 scenes and 6.2 million annotations, 3D-GRAND is the largest 3D-language dataset to date, providing ample data for training and evaluating 3D-LLMs. (2). Dense grounding: Unlike recent million-scale datasets like SceneVerse, which lack grounded language annotations, each language annotation in 3D-GRAND is densely grounded to specific objects or regions within the 3D scenes, facilitating fine-grained language understanding and generation. (3). Diverse language tasks: 3D-GRAND supports a broad array of grounded language tasks, including object reference , spatial reasoning, and scene understanding, making it a comprehensive benchmark for evaluating 3D-LLMs. (4). High-quality annotations: The language annotations in 3D-GRAND are meticulously collected, filtered, and evaluated to ensure they are of high quality, diverse, and natural.
@@
4 3D-POPE: A Benchmark for Evaluating Hallucination in 3D-LLMs
Benchmark design. 3D-POPE consists of a set of triples, each comprising a 3D scene, a posed question, and a binary answer (“Yes” or “No”) indicating the presence or absence of an object (Fig. 1 middle). To ensure a balanced dataset, we maintain a 1:1 ratio of existent to nonexistent objects when constructing these triples. For the selection of negative samples (i.e., nonexistent objects), we employ three distinct sampling strategies:
  • Random Sampling: Nonexistent objects are randomly selected from the set of objects not present in the 3D scene.
  • Popular Sampling: We select the top-$k$ most frequent objects not present in the 3D scene, where $k$ equals the number of objects currently in the scene.
  • Adversarial Sampling: For each positively identified object in the scene, we rank objects that are not present and have not been used as negative samples based on their frequency of co-occurrence with the positive object in the training dataset. The highest-ranking co-occurring object is then selected as the adversarial sample. This approach differs from the original POPE [41] to avoid adversarial samples mirroring popular samples, as indoor scenes often contain similar objects.
These sampling strategies are designed to challenge the model’s robustness and assess its susceptibility to different levels of object hallucination.
Metrics. To evaluate the model’s performance on the 3D-POPE benchmark, we use key metrics including Precision, Recall, F1 Score, Accuracy, and Yes (%). Precision and Recall assess the model’s ability to correctly affirm the presence of objects and identify the absence of objects, respectively. Precision is particularly important as it indicates the proportion of non-existing objects generated by the 3D-LLMs. The F1 Score, combining Precision and Recall, offers a balanced view of performance and serves as the primary evaluation metric. Accuracy measures the proportion of correctly answered questions, encompassing both “Yes” and “No” responses. Additionally, the Yes (%) metric reports the ratio of incorrect “Yes” responses to understand the model’s tendencies regarding object hallucination.
Dataset 3D-POPE Model Precision Recall F1 Score Accuracy Yes (%)
ScanNet200 Val Random Random Baseline 50.00 50.00 50.00 50.00 50.00
3D-LLM [27] 50.03 99.88 66.67 50.07 99.81
3D-VisTA [83] 50.12 53.58 51.79 49.66 53.95
LEO [30] 51.95 77.65 62.25 52.91 74.73
Ours zero-shot (Grounding) 93.34 84.25 88.56 89.12 45.13
Popular Random Baseline 50.00 50.00 50.00 50.00 50.00
3D-LLM [27] 49.97 99.88 66.61 49.94 99.94
3D-VisTA [83] 47.40 51.88 49.54 49.49 52.30
LEO [30] 48.30 77.65 59.55 47.27 80.38
Ours zero-shot (Grounding) 73.05 84.28 78.26 76.59 57.69
Adversarial Random Baseline 50.00 50.00 50.00 50.00 50.00
3D-LLM [27] 49.97 99.88 66.61 49.94 99.94
3D-VisTA [83] 48.28 54.39 51.15 51.14 52.99
LEO [30] 48.47 77.98 59.78 47.52 80.45
Ours zero-shot (Grounding) 69.86 84.21 76.37 73.95 60.26
Model Acc@0.25 Acc@0.5
Non-zero-shot 3D-LLM (flamingo) [27] 21.2 -
3D-LLM (BLIP2-opt) [27] 29.6 -
3D-LLM (BLIP2-flant5) [27] 30.3 -
Zero-shot LLM-Grounder [71] 17.1 5.3
Sim-to-real zero-shot 3D-GRAND 38.0 27.4
Table 3: ScanRefer Results for evaluating visual grounding capability of 3D-LLMs. 3D-GRAND achieves the best zero-shot performance among 3D-LLMs, providing signals for sim-to-real transfer.
3D-POPE Model Precision
Random 3D-GRAND 93.34
w/o grounding tokens 91.96
Popular 3D-GRAND 73.05
w/o grounding tokens 70.37
Adversarial 3D-GRAND 69.86
w/o grounding tokens 67.48

### v3

@@ Abstract
The integration of language and 3D perception is crucial for embodied agents and robots that comprehend and interact with the physical world. While large language models (LLMs) have demonstrated impressive language understanding and generation capabilities, their adaptation to 3D environments (3D-LLMs) remains in its early stages. A primary challenge is a lack of large-scale datasets with dense grounding between language and 3D scenes. We introduce 3D-GRAND, a pioneering large-scale dataset comprising 40,087 household scenes paired with 6.2 million densely-grounded scene-language instructions. Our results show that instruction tuning with 3D-GRAND significantly enhances grounding capabilities and reduces hallucinations in 3D-LLMs. As part of our contributions, we propose a comprehensive benchmark 3D-POPE to systematically evaluate hallucination in 3D-LLMs, enabling fair comparisons of models. Our experiments highlight a scaling effect between dataset size and 3D-LLM performance, emphasizing the importance of large-scale 3D-text datasets for embodied AI research. Our results demonstrate early signals for effective sim-to-real transfer, indicating that models trained on large synthetic data can perform well on real-world 3D scans. Through 3D-GRAND and 3D-POPE, we aim to equip the embodied AI community with resources and insights to lead to more reliable and better-grounded 3D-LLMs.
@@ 1 Introduction
Table 1: Comparison with existing 3D scene datasets with language annotations. 3D-GRAND is the largest language-grounded dataset.
Recent advances in Large Language Models (LLMs) have greatly benefited Embodied Artificial Intelligence (EAI). LLMs demonstrate exceptional capabilities in understanding language instructions [52, 66], perceiving the environment [46, 41, 3, 83, 74], and planning detailed actions [7, 35]. The primary inputs to LLMs, other than language, have been the combination of language and 2D images, categorizing these models as 2D-LLMs. The significant advancements in 2D-LLMs can be mainly attributed to their training on extensive vision-language datasets. These datasets [60, 84], with billions of image and text pairs, have been instrumental in enhancing the models’ understanding of visual content and its contextual relevance to text information. These datasets have provided the foundational data needed for training models excelling at integrating vision and language. Despite some progress in equipping LLMs to understand 3D scenes (3D-LLMs) [28, 30, 68, 31, 85, 11, 55], these models are limited by the scarcity of 3D scene and text pairs. In this work, we introduce 3D-GRAND, a pioneering million-scale dataset designed for densely-grounded 3D Instruction Tuning.
Recently, SceneVerse [37] concurrently introduced a large-scale 3D vision-language dataset. However, a significant limitation of this dataset is the absence of object grounding in language, which is crucial for usability in robotics tasks and reducing hallucination. Research on 2D-LLMs indicates that grounding language to 2D contexts mitigates hallucination in language models [76, 53, 5, 40, 56, 80], thus enhancing the reliability and interpretability of generated responses. While 2D grounding has been extensively explored, extending these principles to 3D environments is still underdeveloped. This situation raises two critical questions: (1) Is there hallucination in 3D-LLMs and if so, how severe is it? (2) Can densely-grounded data mitigate hallucination for 3D-LLMs? These questions underscore a critical need within the research community for the development of an evaluation benchmark specifically designed for 3D-LLMs and the construction of a large-scale, 3D-grounded dataset.
To quantify hallucination in 3D LLMs, this work introduces 3D-POPE (3D Polling-based Object Probing Evaluation). 3D-POPE provides a comprehensive and standardized protocol for evaluating hallucination that enables systematic assessment and facilitates fair comparisons across 3D-LLMs, enhancing our understanding of model capabilities in object hallucination. Specifically, we pose existence questions to 3D-LLMs and evaluate their responses, as shown in Fig LABEL:fig:main-teaser.
To evaluate the role of dense-grounding, we introduce a pioneering million-scale dataset, 3D-GRAND, for densely grounded 3D instruction tuning. 3D-GRAND includes 40K household scenes paired with 6.2 million scene-language instructions, featuring dense phrase-to-object grounding. We conduct rigorous human evaluations to ensure the dataset’s quality. Our results trained with 3D-GRAND highlight the dataset’s effectiveness in enhancing grounding and reducing hallucination for 3D-LLMs. We highlight the effectiveness of incorporating 3D-GRAND in Fig LABEL:fig:main-teaser and introduce each category and provide examples in Fig 1.
Our contributions include: (1) 3D-GRAND, the first million-scale, densely-grounded 3D-text dataset for grounded 3D Instruction Tuning. 3D-GRAND includes 40K household scenes paired with 6.2M densely-grounded scene-language instructions. (2) 3D-POPE, a suite of benchmarks and metrics that systematically evaluate hallucination, enabling fair comparisons of future 3D-LLM models in terms of object hallucination. and (3) Findings for hallucination, grounding, and scaling that guide future research, including: (a) training with 3D-GRAND significantly reduces hallucinations particularly when the data is densely grounded; (b) densely grounded instruction tuning significantly enhances the grounding capabilities of 3D-LLMs; (c) scaling densely grounded data consistently improves grounding accuracy and reduces hallucination; and (d) models transfer from sim-to-real, providing an early signal for a low-cost and sustainable future of scaling synthetic 3D data to help on real tasks.
@@ 2 Related Work
Injecting 3D into LLMs. Recent advances in large language models (LLMs) have inspired work extending their capabilities to 3D environments, leading to the development of 3D-LLMs [11, 55, 74, 85]. Notable works include 3D-LLM [28], which integrates 3D point clouds and features into LLMs to enable tasks like captioning, question answering, and navigation. LEO [31] excels as an embodied multi-modal generalist agent in perception, grounding, reasoning, planning, and action in 3D environments, showing the potential of 3D-LLMs in understanding and interacting with the physical world. The most relevant work is Chat-3Dv2 [68, 30], which grounds generated scene captions to objects in 3D scenes. However, Chat-3Dv2’s dataset is limited to one type of 3D-text task (scene captioning) and only has 705 captions from a subset of ScanNet scenes. In 3D-GRAND, we expand this concept by diversifying 3D-text tasks and increasing data scale to million-scale. Our results show promising data scaling effects and sim-to-real transfer, paving the way for future large-scale 3D-LLM training.
@@
Object Hallucination of VLMs. While 2D VLMs have achieved impressive performance, they are prone to hallucinating objects that do not exist in the provided images, a problem known as object hallucination [58, 16, 12]. Several methods have been suggested to mitigate the object hallucination issue, such as integrating an external object detector [79], applying visually grounded visual instruction tuning [76, 80] or reinforcement learning [63, 25], performing iterative refinement [82], and adapting the decoding strategies [33]. To quantify and mitigate this issue, several benchmarks have been proposed. CHAIR (Caption Hallucination Assessment with Image Relevance) [58] measures the frequency of hallucinated objects in image captions by comparing the objects mentioned to the ground truth annotations. POPE (Probing Object Hallucination Evaluation) [43] assesses a VLM’s ability to identify the presence or absence of objects through yes/no probing questions. However, these studies primarily focus on 2D image-text datasets like COCO [45]. In contrast, object hallucination in 3D-LLMs remains largely unexplored. Our work addresses this gap by introducing 3D-POPE, a comprehensive benchmark for evaluating object hallucination in 3D-LLMs. To the best of our knowledge, this is the first object hallucination benchmark for 3D-LLMs.
@@
Grounding Datasets for 3D-LLMs. In 2D, large-scale datasets with grounding information have been instrumental in vision-language research. Notable examples include RefCOCO [77], which provides referring expressions for objects in COCO images [45]. Additionally, 2D LLMs [53, 56, 71, 40, 76] have been trained with densely-grounded web-crawled image-text pairs. In 3D, there is a growing interest in creating datasets that pair 3D scenes with textual annotations [78, 1, 34, 13]. ScanRefer [10] pioneered this effort by contributing a dataset of ScanNet [15] scenes with referring expressions. Table 1 introduces the efforts made by the community. However, these datasets have limited grounding annotations and often focus on a single task, such as referring expression comprehension or visual question answering. In contrast, our proposed dataset, 3D-GRAND, stands out by providing 6.2 million densely-grounded scene-language instructions across a diverse set of 3D-text tasks and 40,087 household scenes. This enables a wide range of grounding tasks and facilitates the development of more reliable and better-grounded 3D-LLMs.
Of recent datasets, SceneVerse [34] is the most similar to ours: both are million-scale 3D grounding datasets. A key difference is that SceneVerse [34] provides sparse grounding, while 3D-GRAND is densely grounded. For instance, Sceneverse grounds “This is a big cotton sofa. It is between the window and the wooden table.” to a single entity (e.g., [sofa-3]). Instead, 3D-GRAND grounds every noun phrase in the caption: “big cotton sofa”, “it”, “window” and “wooden table” to an entity.
For clarity of describing the differences between approaches and datasets, we define the grounding granularity as follows and show an example in Appendix A: Paragraph-level set-to-set grounding: Many sentences in a long paragraph, each containing several object nouns, are linked to a set of 3D objects without clear associations from specific sentences/noun phrases to objects. Session-level many-to-one grounding: Multiple sentences in one session, where each can describe several objects (targets and landmarks), are associated with one 3D object. Noun-level one-to-one grounding: Each noun phrase in each sentence is explicitly matched with one 3D object.
Scene Caption Object Reference QA
SceneVerse Paragraph-level set-to-set grounding Session-level many-to-one grounding No grounding
3D-GRAND Noun-level one-to-one grounding Noun-level one-to-one grounding Noun-level one-to-one grounding
Table 2: Comparison of grounding granularity in SceneVerse and 3D-GRAND.
Additionally, the language annotations of 3D-GRAND are more trustworthy and have higher quality. Hallucination is known as one of the most common mistakes of LLMs [32, 43, 58]. In 3D-GRAND, we use a hallucination filter to check and delete any annotations with hallucinated object IDs. This is not possible for SceneVerse since they have pure language output. 3D-GRAND is also quality-checked by humans.
3 3D-POPE: A Benchmark for Evaluating Hallucination in 3D-LLMs
3D-POPE Model Precision Recall F1 Score Accuracy Yes (%)
Random Random Baseline 50.00 50.00 50.00 50.00 50.00
3D-LLM [28] 50.03 99.88 66.67 50.07 99.81
3D-VisTA [85] 50.12 53.58 51.79 49.66 53.95
LEO [31] 51.95 77.65 62.25 52.91 74.73
Ours 0-shot (Grounding) 93.34 84.25 88.56 89.12 45.13
Popular Random Baseline 50.00 50.00 50.00 50.00 50.00
3D-LLM [28] 49.97 99.88 66.61 49.94 99.94
3D-VisTA [85] 47.40 51.88 49.54 49.49 52.30
LEO [31] 48.30 77.65 59.55 47.27 80.38
Ours 0-shot (Grounding) 73.05 84.28 78.26 76.59 57.69
Adversarial Random Baseline 50.00 50.00 50.00 50.00 50.00
3D-LLM [28] 49.97 99.88 66.61 49.94 99.94
3D-VisTA [85] 48.28 54.39 51.15 51.14 52.99
LEO [31] 48.47 77.98 59.78 47.52 80.45
Ours 0-shot (Grounding) 69.86 84.21 76.37 73.95 60.26
Benchmark design. 3D-POPE consists of triples with: a 3D scene, a posed question, and a binary answer (“Yes” or “No”) indicating the presence or absence of an object (Fig. LABEL:fig:main-teaser middle). For balance, we maintain a 1:1 ratio of existent to nonexistent objects when constructing these triples. For negative samples (i.e., nonexistent objects), we use the following three distinct sampling strategies that are designed to challenge the model’s robustness and assess its susceptibility to different levels of object hallucination. Random Sampling: Nonexistent objects are randomly selected from the set of objects not present in the 3D scene. Popular Sampling: We select the top-$k$ most frequent objects not present in the 3D scene, where $k$ equals the number of objects currently in the scene. Adversarial Sampling: For each positively identified object in the scene, we rank objects that are not present and have not been used as negative samples based on their frequency of co-occurrence with the positive object in the training dataset. The highest-ranking co-occurring object is selected as the adversarial sample, which differs from the original POPE [43] to avoid adversarial samples mirroring popular samples, as indoor scenes often contain similar objects.
Metrics. 3D-POPE metrics include Precision, Recall, F1 Score, Accuracy, and Yes (%). Precision and Recall assess the model’s ability to correctly affirm the presence of objects and identify the absence of objects, respectively. Precision is particularly important as it indicates the proportion of non-existing objects generated by the 3D-LLMs. The F1 Score, combining Precision and Recall, balances the two and serves as the primary evaluation metric. Accuracy measures the proportion of correctly answered questions for “Yes” and “No” responses. Additionally, the Yes (%) metric reports the ratio of incorrect “Yes” responses to understand the model’s tendencies regarding object hallucination.
@@
4 3D-GRAND: 3D Ground Anything Dataset
We introduce 3D-GRAND, a large-scale, densely-grounded 3D-text dataset designed for grounded 3D instruction tuning. We describe the data collection process, dataset statistics, and the unique features that make 3D-GRAND a valuable resource for advancing research in 3D-LLMs.
@@
Thus, we curate a diverse collection of 40,087 high-quality 3D indoor scenes from the 3D-FRONT [24] and Structured3D [81] datasets. These datasets are chosen for their large quantities of synthetic indoor scenes with professionally designed layouts. The collection includes various room types, like living rooms, bedrooms, kitchens, office spaces, and conference rooms. We process these 3D scenes to generate per-room 3D point clouds. Details on point cloud rendering and cleaning are provided in the Appendix.
@@
The second, and perhaps more critical, motivation is the scalability of annotation. While we can potentially scale up 3D scenes using synthetic data generation pipelines, manual annotation is both costly and time-consuming, especially for complex tasks like densely grounding annotation. To put the time/money cost in perspective, for the data we annotated in this paper, we estimate that obtaining the same annotations with human annotator would cost at least \$539,000 and require 5.76 years (no eat, no sleep) worth of work from a professional annotator (earning minimum wage of \$10.67 per hour). In contrast, using LLMs (GPT4 [52]), we achieve the same results for \$3,030 within 2 days, representing a 178x reduction in cost and a 1051x reduction in time. At the time of writing, the cost and time further decreases by 50% to \$1,500 and 1 day, with the introduction of GPT-4o [51].
@@
$\bullet$ 3D Model to 2D Image. In the 3D-Front dataset, each object is sourced from 3D-Future [24], which provides a ground truth 2D image for each object. For the Structured3D dataset, individual images for each object are not available. Thus, we use set-of-mark prompting [73], where each object to be annotated is circled in red in the images.
$\bullet$ 2D Image to Attributes. We use GPT-4V to generate detailed language annotations for each 2D object image, including attributes like name, color, finish, and texture. The naming is open-vocabulary, contrary to being class-agnostic.
@@
$\bullet$ Scene Graph to Generated Annotations. Based on the given scene graph, we will be able to produce 3D-Grounded Object Reference, 3D-Grounded Scene Description, and 3D-Grounded QA using GPT-4 [52] with various prompts, which we will show in the appendix.
$\bullet$ Generated Annotations to Processed Annotations. After we acquire raw annotations, we will apply hallucination filters and template augmentation for the phrase tag to remove low-quality annotations and augment generated annotations.
With this pipeline, we generate a diverse range of 3D vision-language understanding tasks as shown in Figure 1. On a high level, these tasks can be categorized into:
@@
Annotation Source Error Rate
ScanEnts3D (AMT) 16%
ScanEnts3D (Professional) $<$5%
3D-GRAND (LLM, GPT-4) 5.6-8.2%
Table 4: Error rates comparison between ScanEnts3D and 3D-GRAND annotations. (AMT = Amazon Mechanical Turk)
Human quality check. Table 4 shows results from extensive human quality checks on 5,100 generated annotations, comparing the error rates of 3D-GRAND and previous datasets such as ScanEnts3D [1]. The results show that large language models (LLMs) like GPT-4 can achieve error rates in densely grounded annotations comparable to those of professional human annotators. This finding aligns with studies that suggest LLMs are starting to reach human-level annotation quality on certain tasks [65]. See supplement for a more detailed description of the human quality check process.
Dataset highlights. 3D-GRAND has several features that distinguish it from existing 3D-language datasets: (1). Large-scale: With 40,087 scenes and 6.2 million annotations, 3D-GRAND is the largest 3D-language dataset to date, providing ample data for training and evaluating 3D-LLMs. (2). Dense grounding: Unlike recent million-scale datasets like SceneVerse, which lack grounded language annotations, each language annotation in 3D-GRAND is densely grounded to specific objects or regions within the 3D scenes, facilitating fine-grained language understanding and generation. (3). Diverse language tasks: 3D-GRAND supports many grounded language tasks, including object reference , spatial reasoning, and scene understanding, making it a comprehensive benchmark for evaluating 3D-LLMs. (4). High-quality annotations: We utilize a hallucination filter to mitigate hallucination of the language annotations in 3D-GRAND. They are also human-evaluated to ensure the quality.